Merry is selling her old car for $13252. If a prospective buyer offers to make yearly payments for the next 5 years, what should Merry set these payments to if the correct discount rate is 0.21?
Answer #1

This question requires application of PV of annuity formula

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

13252 =P*1-(1+0.21)-5 13252 =P* 0.21

13252 = P * 2.925984

P = $4,529.07